Determine the Integrating factor and identify the solution of general/initial value problem

a) (π‘₯2 +1)𝑦′ +2π‘₯𝑦=4π‘₯2 ; 𝑦(0)=0

1+π‘₯2


b)𝑦′+ 𝑦 =3π‘₯2 π‘₯𝑙𝑛π‘₯ ln π‘₯

c) (π‘₯2 + 1) 𝑦′ + 4π‘₯𝑦 = π‘₯ ; 𝑦(2) = 1


1
Expert's answer
2021-05-25T13:14:02-0400

(a) Given equation-

"(x^2+1)y'+2xy=4x^2; y(0)=0"


The simplifying expression is-


"\dfrac{dy}{dx}+\dfrac{2xy}{x^2+1}=\dfrac{4x^2}{x^2+1}"



Integrating factor-


"I.F.=e^{\int Pdx}"

Β Β "=e^{\frac{2x}{1+x^2}}dx



 \\=e^{ln(1+x^2)}
\\
 =1+x^2"

Solution is-


"y\times (1+x^2)=\int (1+x^2)(\dfrac{4x^2}{1+x^2})dx

\\[9pt]

\Rightarrow y(1+x^2)=\dfrac{4x^3}{3}+C"


Putting Y(0)=0


"0=0+C\Rightarrow C=0"


So Solution is-


"(1+x^2)y=\dfrac{4x^3}{3}"


(b)Given equation-

Β Β "y'+y=3x^2
\\[9pt]


 \dfrac{dy}{dx}+y=3x^2"


Integrating factor-


I".F.=e^{1.dx}=e^x"


Solution is-


"y\times e^x=\int e^x 3x^2dx


\\[9pt]
ye^x=3x^2e^x-6xe^x+6e^x+C
\\[9pt]


y=3x^2-6x+6+ce^{-x}"


(C) Given equation is-


"(x^2+1)y'+4xy=x;y(2)=1"


The simplified expression is-


"\dfrac{dy}{dx}+\dfrac{4xy}{x^2+1}=\dfrac{x}{x^2+1}"



Integrating factor-


"I.F.=e^{\frac{4x}{1+x^2}dx}=e^{2log(1+x^2)}=e^{log(1+x^2)^2}=(1+x^2)^2"


Solution is-


"y\times x^2=\int (1+x^2)^2 \dfrac{x}{x^2+1}dx
\\[9pt]

\Rightarrow yx^2=\int(x+x^3)dx\\[9pt]



\Rightarrow yx^2=\dfrac{x^2}{2}+\dfrac{x^4}{4}+c


\\[9pt]
\Rightarrow y=\dfrac{1}{2}+\dfrac{x^2}{4}+\dfrac{c}{x^2}"


Also, y(2)=1


"\Rightarrow 1=\dfrac{1}{2}+\dfrac{4}{4}+\dfrac{c}{4}\\[9pt]



\dfrac{c}{4}=-\dfrac{1}{2}\\[9pt]c=-2"


So solution is-

"y=\dfrac{1}{2}+\dfrac{x^2}{4}-\dfrac{2}{x^2}"